# Alleviating the Hubble tension with Torsion Condensation (TorC)

### v1

@@ II.3 Cosmological effects of $\Omega_{L}$ and $\varpi_{r}$
The effects of $\varpi_{r}$ on the Hubble horizon and CMB power spectrum are shown in Fig. 3. The left panel shows that a decreasing $\varpi_{r}$ reduces the comoving Hubble horizon in the early Universe, while converging with $\Lambda\mathrm{CDM}$ at late times. Correspondingly, the right panel shows that a smaller $\varpi_{r}$ shifts the acoustic peaks to higher multipoles $\ell$ and increases their amplitudes. This shift results from the reduced sound horizon at decoupling, which is a consequence of the smaller Hubble horizon at early times. This mechanism is phenomenologically similar to that of early dark energy models (EDE) [38, 39]. However, unlike many EDE models, which introduce a new scalar field and potential for phenomenological gain, the torsional field in TorC arises uniquely from the unitary and power-counting renormalizable gauging of the Poincaré group, providing a more fundamental motivation. Furthermore, the smaller early Hubble horizon intensifies gravitational potential decay during the radiation-dominated era, thereby enhancing the early ISW effect, which acts as an additional driving force for photon-baryon oscillations and boosts the peak amplitudes.
@@ III   Methods
The previous sections introduced the two new parameters in TorC, $\Omega_{L}$ and $\varpi_{r}$, and their effects on the Hubble horizon and CMB power spectra of the TorC Universe. To find the posterior distribution of the cosmological parameters in TorC, the Polychord [40] nested sampling algorithm is employed to explore the parameter space of TorC cosmological parameters. The sampling is conducted using Cobaya [32] which interfaces the CAMB Boltzmann solver with Polychord and relevant cosmological likelihoods. Note that this analysis modifies the background expansion in CAMB using the TorC effective dark energy density and pressure, while the perturbation equations remain those of standard $\Lambda\mathrm{CDM}$ [^8] As a initial study, only the background evolution of TorC is being investigated, cosmological perturbation theory for TorC will be developed in future work.
@@ III.1 Numerical implementation
As discussed in Section II.3, to explore the full parameter space of $\Omega_{L}$ and $\varpi_{r}$, the CAMB Boltzmann code is modified to read in a tabulated input of dark energy density $\rho_{\Lambda}(a)$, and pressure $P_{\Lambda}(a)$ as functions of scale factor $a$. The modified version is available at [37]. This customized CAMB is then linked to a modified Cobaya which reads the same external $\rho_{\Lambda}(a)$ and $P_{\Lambda}(a)$, available at [41]. The chains and supplementary files used to produce the plots and results presented in this paper are available at [42].
@@ III.2 Prior and likelihoods
$\Omega_{L}$ $[0.1,1.5]$
$\varpi_{r}$ $[0.1,1.5]$
@@
$\Omega_{b}h^{2}$ $[0.005,0.1]$
$\Omega_{c}h^{2}$ $[0.005,0.99]$
@@ III.3 Bayesian model comparison
$\log\mathcal{Z}=\langle\log\mathcal{L}(\theta)\rangle-D_{KL}.$ (19)
@@
$\langle\log\mathcal{L}(\theta)\rangle=\int\log\mathcal{L}(\theta)P(\theta|D)\mathrm{d}\theta,$ (20)
@@
$D_{KL}=\int P(\theta|D)\log\frac{P(\theta|D)}{\pi(\theta)}\mathrm{d}\theta,$ (21)
which quantifies how much the posterior distribution has contracted relative to the prior. This term acts as a complexity penalty – a model with a large parameter space that is significantly narrowed by the data will have a larger value of $D_{KL}$.
@@
$\begin{split}\Delta\log\mathcal{Z}&=\log\mathcal{Z}_{1}-\log\mathcal{Z}_{2}.\\ &=\Delta\langle\log\mathcal{L}(\theta)\rangle-\Delta D_{KL},\end{split}$ (23)
values of $\Delta\log\mathcal{Z}>0$ indicate that $M_{1}$ is supported over $M_{2}$ by a factor of $\exp(\Delta\log\mathcal{Z})$, while $\Delta\log\mathcal{Z}<0$ suggests the opposite. As shown in Eq. 23, this framework allows the source of model preference to be identified, comparing the goodness-of-fit and the complexity penalty of the two models. This analysis will be performed between TorC and $\Lambda\mathrm{CDM}$, exploring the added complexity of the two new parameters $\Omega_{L}$ and $\varpi_{r}$ in TorC, and their impact on the fit to the datasets.
@@ IV   Results
Figure 4 shows the full posterior distributions of the TorC and $\Lambda\mathrm{CDM}$ cosmological parameters. The result is obtained from PolyChord nested sampling with 1000 live points.
@@
Figure 5: The left panel shows $\Omega_{m}$-$H_{0}$ posterior contours for $\Lambda\mathrm{CDM}$ (using Planck 2018 data [Purple], TRGB data [Orange] [48], and SH0ES 2020 data [Blue] [29]) and for TorC (using Planck 2018 data [Yellow], and joint Planck and SH0ES data [Red]). The contours show the tension between the Planck and SH0ES posteriors in the $\Lambda\mathrm{CDM}$ model, and the alleviation of this tension in the TorC model. The right panel shows the posterior distributions of the two TorC parameters $\Omega_{L}$ and $\varpi_{r}$ in the $H_{0}$-plane. The plot shows a positive correlation between $H_{0}$ and $\Omega_{L}$, and a negative correlation between $H_{0}$ and $\varpi_{r}$.
@@ IV.1 $H_{0}$ Tension
The TorC model is capable of alleviating the $H_{0}$ tension by allowing for a higher inferred Hubble parameter compared to the $\Lambda\mathrm{CDM}$ model. This is shown in the $H_{0}$-$\Omega_{m}$ plane on the left panel of Fig. 5. The $\Lambda\mathrm{CDM}$ posteriors with Planck, SH0ES, and tip of the red giant branch (TRGB) [48] likelihoods reveal a clear tension between the Planck and SH0ES datasets. Whereas the TorC posteriors for both Planck likelihood alone and joint SH0ES and Planck likelihoods show a fit that is consistent with both datasets.
The right panel of Fig. 5 shows the posterior distributions of the two TorC parameters, $\Omega_{L}$ and $\varpi_{r}$, in the $H_{0}$-plane. The $H_{0}$-$\varpi_{r}$ plane shows a negative correlation: as $\varpi_{r}$ decreases, the value of $H_{0}$ increases. This behavior suggests that while the TorC model is consistent with $\Lambda\mathrm{CDM}$, it allows for a higher value of $H_{0}$ through the presence of torsion. Furthermore, the $H_{0}$-$\Omega_{L}$ plane shows a positive correlation, indicating that a higher $H_{0}$ value in TorC is achieved not only through a reduction in $\varpi_{r}$, but also by an increase in the amount of bare dark energy $\Omega_{L}$.
@@ IV.3 Model comparison
Writing the logarithmic Bayes factor in the form of the second line of Eq. 23, namely $\log\Delta\mathcal{Z}=\Delta\langle\log\mathcal{L}(\theta)\rangle-\Delta D_{KL}$, the result with Planck 2018 alone is
@@ V   Conclusions
Future work will extend the analysis of TorC parameters by incorporating other cosmological likelihoods, such as the recent DESI BAO dataset [51]. The present study assumes a vanishing torsional dark energy component, $\lambda=0$, future investigations will consider scenarios in which both bare and torsional contributions to dark energy are present. The inclusion of torsional dark energy may offer a natural explanation for the remarkably small observed value of the cosmological constant. This study is also conducted in the context of a flat Universe. Future work will also explore scenarios with non-zero spatial curvature. While the TorC field equations for background evolution are insensitive to the spatial curvature constant $k$, spatial curvature can still manifest through its standard geometric effects on light propagation and large-scale structure. It will be investigated whether TorC under non-zero spatial curvature can alleviate the curvature tension observed in the $\Lambda\mathrm{CDM}$ model. Finally, the development of a perturbation theory for TorC will enable a deeper exploration of the model impact on cosmic structure formation and the CMB.
@@ Acknowledgements.
The authors would like to thank Mike Hobson, Anthony Lasenby, Adam Ormondroyd, and Dily Ong for useful discussions. SL would like to thank the IBM colour blind friendly palette for the color scheme used in the figures, and Joe Pattison and Oscar O’Hara for their advice on making the plots more visually appealing. WB is grateful for the support of Girton College, Cambridge, Marie Skłodowska-Curie Actions and the Institute of Physics of the Czech Academy of Sciences. This work used the DiRAC Data Intensive service (CSD3 www.csd3.cam.ac.uk) at the University of Cambridge, managed by the University of Cambridge University Information Services on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC component of CSD3 at Cambridge was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. The authors acknowledge the use of NumPy [52], SciPy [53], and Matplotlib [54] for numerical computations and plotting, and gemini-2.5-flash-preview-04-17 for the improvement of clarity of this paper.
@@ Appendix A Condensation and perturbation
$\tensor{\mathcal{R}}{{}^{\rho}_{\sigma}{}_{\mu\nu}}$ $\equiv 2\left(\tensor{\partial}{{}_{[\mu|}}\tensor{\Gamma}{{}^{\rho}_{|\nu]}{}_{\sigma}}+\tensor{\Gamma}{{}^{\rho}_{[\mu|}{}_{\alpha}}\tensor{\Gamma}{{}^{\alpha}_{|\nu]}{}_{\sigma}}\right)\,,$ (28)
@@
$\tensor{\Gamma}{{}^{\rho}_{\mu}{}_{\nu}}\equiv\big{\{}\tensor{}{{}^{\rho}_{\mu}{}_{\nu}}\big{\}}+\frac{1}{2}\left(\tensor{\mathcal{T}}{{}^{\rho}_{\mu\nu}}-\tensor{\mathcal{T}}{{}_{\nu\mu}^{\rho}}+\tensor{\mathcal{T}}{{}_{\mu}^{\rho}{}_{\nu}}\right),$ (30)
@@
where $\tensor{\nabla}{}$ denotes the usual GR covariant derivative with connection {}$\big{\{}\tensor{}{{}^{\nu}_{\mu}{}_{\rho}}\big{\}}$ and where the Riemannian curvature tensor of GR is
@@
Figure 7: Output generated by PSALTer. The spectrograph of the theory defined by Eq. 1. At the top, the (Fourier transformed) wave operator $\mathsf{O}(k)$ from Eq. 34 is displayed in the $J^{P}$-representation defined in Tables 4 and 5. The $J^{P}$ states are organised into diagonal blocks of $J=0$$J=1$ and $J=2$, and within each block into diagonal parity-preserving sectors of $P=1$ (orange) and $P=-1$ (purple), along with off-diagonal parity-violating blocks (teal). The Lagrangian in Eq. 1 does not contain any odd powers of the totally antisymmetric tensor, so there is no mixing between parities in this theory. The theory contains as many as 21 gauge generators: this far exceeds the 10 gauge generators common to all instances of Poincaré gauge theory. The gauge symmetries impose constraints on the source currents. The particle spectrum consists of two massless polarisations and one massive pseudoscalar particle (i.e. $J^{P}=0^{-}$). In terms of the couplings in Eq. 1 the pseudoscalar has mass $\sqrt{\lambda}$, so that it is not a tachyon if $\lambda\geq 0$. Its pole residue indicates that it is not a ghost if $\mu<0$. The residue of the massless pole (where $p$ denotes the particle energy in some fiducial frame) then indicates a no-ghost condition of $(\nu+2\mu)(\nu-\mu)>0$. Together, these three unitarity conditions are consistent with those in Eq. 2.
@@ Appendix B Initial conditions
$a=$ $\frac{H_{0}\sqrt{\Omega_{r}}}{\alpha\varpi_{r}}\tau+\frac{H_{0}^{2}\Omega_{m}\bigl{(}3\varpi_{r}^{2}+1\bigl{)}}{16\alpha\varpi_{r}^{2}}\tau^{2}$
$+\frac{5H_{0}^{3}\Omega_{m}^{2}\bigl{(}\varpi_{r}^{2}-1\bigr{)}}{512\alpha\varpi_{r}^{3}\sqrt{\Omega_{r}}}\tau^{3}$
$+\frac{H_{0}^{4}\bigl{(}\varpi_{r}^{2}-1\bigr{)}\bigl{(}27\varpi_{r}^{2}-121\bigr{)}\Omega_{m}^{3}}{49152\alpha\varpi_{r}^{4}\Omega_{r}}\tau^{4}$
$+\frac{H_{0}^{5}}{131720\alpha\varpi_{r}^{5}\Omega_{r}^{3/2}}\biggl{(}32768\bigl{(}3\varpi_{r}^{2}+1\bigr{)}\Omega_{r}^{3}\Omega_{L}$
$-49\bigl{(}9\varpi_{r}^{4}-29\varpi_{r}^{2}+20\bigr{)}\Omega_{m}^{4}\biggr{)}\tau^{5}+\mathcal{O}(\tau^{6}),$ (36a)
ϖ=italic-$\varpi=$ $\varpi_{r}-\frac{3H_{0}\bigl{(}\varpi_{r}^{2}-1\bigr{)}\Omega_{m}}{16\sqrt{\Omega_{r}}}\tau$
$+\frac{H_{0}^{2}\bigl{(}18\varpi_{r}^{4}-5\varpi_{r}^{2}-13\bigr{)}\Omega_{m}^{2}}{512\varpi_{r}\Omega_{r}}\tau^{2}$
$+\frac{H_{0}^{3}\bigl{(}-324\varpi_{r}^{6}+45\varpi_{r}^{4}-20\varpi_{r}^{2}+299\bigr{)}\Omega_{m}^{3}}{49152\varpi_{r}^{2}\Omega_{r}^{3/2}}\tau^{3}$
$+\frac{H_{0}^{4}}{1310720\varpi_{r}^{3}\Omega_{r}^{2}}\biggl{[}-98304\Omega_{r}^{3}\Omega_{L}$
$+\bigl{(}\varpi_{r}^{2}-1\bigr{)}\biggl{(}\bigl{(}2\varpi_{r}^{2}\bigl{(}810(\varpi_{r}^{2}+\varpi_{r}^{4})+731\bigr{)}$
$+2327\bigr{)}\biggr{)}\Omega_{m}^{4}\biggr{]}\tau^{4}+\mathcal{O}(\tau^{5}).$ (36b)
The parameter $\varpi_{r}$ corresponds to the zeroth-order term in the power series of $\varpi$. As such, the power series of $a$ and $\varpi$ are functions of $\varpi_{r}$ and $\tau$. Note that the normalisation $\alpha$ only affects the scale factor $a$. To ensure the validity of the power series, the initial value of $\tau$ used for numerical evolution is chosen to be sufficiently small such that the first term in the power series of $a$ dominates over the second term
$\tau_{\mathrm{ini}}=10^{-6}\frac{16\sqrt{\Omega_{r}}\varpi_{r}}{\Omega_{m}\bigl{(}3\varpi_{r}^{2}+1\bigr{)}H_{0}}.$ (37)
At such an early time, it is clear that the system’s evolution is primarily governed by the zeroth-order term in the power series expansion of the torsion scalar field, $\varpi_{r}$.

### v4

@@ II.3 Cosmological effects of $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$
The effects of $\varpi_{\mathrm{r}}$ on the Hubble horizon and CMB power spectrum are shown in Fig. 3. The left panel shows that a decreasing $\varpi_{\mathrm{r}}$ reduces the comoving Hubble horizon in the early Universe, while converging with $\Lambda\mathrm{CDM}$ at late times. Correspondingly, the right panel shows that a smaller $\varpi_{\mathrm{r}}$ shifts the acoustic peaks to higher multipoles $\ell$ and increases their amplitudes. This shift results from the reduced sound horizon at decoupling, which is a consequence of the smaller Hubble horizon at early times. This mechanism is phenomenologically similar to that of early dark energy models (EDE) [42, 43]. However, unlike many EDE models, which introduce a new scalar field and potential for phenomenological gain, the torsional field in TorC arises uniquely from the unitary and power-counting renormalizable gauging of the Poincaré group, providing a more fundamental motivation. Furthermore, the smaller early Hubble horizon intensifies gravitational potential decay during the radiation-dominated era, thereby enhancing the early ISW effect, which acts as an additional driving force for photon-baryon oscillations and boosts the peak amplitudes.
@@ III  Methods
The previous sections introduced the two new parameters in TorC, $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$, and their effects on the Hubble horizon and CMB power spectra of the TorC Universe. To find the posterior distribution of the cosmological parameters in TorC, the Polychord [44] nested sampling algorithm is employed to explore the parameter space of TorC cosmological parameters. The sampling is conducted using Cobaya [36] which interfaces the CAMB Boltzmann solver with Polychord and relevant cosmological likelihoods. Note that this analysis modifies the background expansion in CAMB using the TorC effective dark energy density and pressure, while the perturbation equations remain those of standard $\Lambda\mathrm{CDM}$. [^8] As a initial study, only the background evolution of TorC is being investigated, cosmological perturbation theory for TorC will be developed in future work.
@@ III.1 Numerical implementation
As discussed in Section II.3, to explore the full parameter space of $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$, the CAMB Boltzmann code is modified to read in a tabulated input of dark energy density $\rho_{\Lambda}(a)$, and pressure $P_{\Lambda}(a)$ as functions of scale factor $a$. The modified version is available at [41]. This customized CAMB is then linked to a modified Cobaya which reads the same external $\rho_{\Lambda}(a)$ and $P_{\Lambda}(a)$, available at [45]. The chains and supplementary files used to produce the plots and results presented in this paper are available at [46].
@@ III.2 Prior and likelihoods
$\Omega_{\mathrm{L}}$ $[0.1,1.5]$
$\varpi_{\mathrm{r}}$ $[0.1,1.5]$
@@
$\Omega_{\mathrm{b}}h^{2}$ $[0.005,0.1]$
$\Omega_{\mathrm{c}}h^{2}$ $[0.005,0.99]$
@@ III.3 Bayesian model comparison
$\log\mathcal{Z}=\langle\log\mathcal{L}(\theta)\rangle_{\mathcal{P}}-\mathcal{D}_{\mathrm{KL}}.$ (19)
@@
$\langle\log\mathcal{L}(\theta)\rangle_{\mathcal{P}}=\int\log\mathcal{L}(\theta)P(\theta|D)\mathrm{d}\theta,$ (20)
@@
$\mathcal{D}_{\mathrm{KL}}=\int P(\theta|D)\log\frac{P(\theta|D)}{\pi(\theta)}\mathrm{d}\theta,$ (21)
which quantifies how much the posterior distribution has contracted relative to the prior. This term acts as a complexity penalty – a model with a large parameter space that is significantly narrowed by the data will have a larger value of $\mathcal{D}_{\mathrm{KL}}$.
@@
$\begin{split}\Delta\log\mathcal{Z}&=\log\mathcal{Z}_{1}-\log\mathcal{Z}_{2}.\\ &=\Delta\langle\log\mathcal{L}(\theta)\rangle_{\mathcal{P}}-\Delta\mathcal{D}_{\mathrm{KL}},\end{split}$ (23)
values of $\Delta\log\mathcal{Z}>0$ indicate that $M_{1}$ is supported over $M_{2}$ by a factor of $\exp(\Delta\log\mathcal{Z})$, while $\Delta\log\mathcal{Z}<0$ suggests the opposite. As shown in Eq. 23, this framework allows the source of model preference to be identified, comparing the goodness-of-fit and the complexity penalty of the two models. This analysis will be performed between TorC and $\Lambda\mathrm{CDM}$, exploring the added complexity of the two new parameters $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$ in TorC, and their impact on the fit to the datasets.
@@ IV  Results
Fig. 4 shows the full posterior distributions of the TorC and $\Lambda\mathrm{CDM}$ cosmological parameters. The result is obtained from PolyChord nested sampling with 1000 live points.
@@
Figure 5: The left panel shows $\Omega_{\mathrm{m}}$-$H_{0}$ posterior contours for $\Lambda\mathrm{CDM}$ (using Planck 2018 data [Purple], TRGB data [Orange] [52], and SH0ES 2020 data [Blue] [33]) and for TorC (using Planck 2018 data [Yellow], and joint Planck and SH0ES data [Red]). The contours show the tension between the Planck and SH0ES posteriors in the $\Lambda\mathrm{CDM}$ model, and the alleviation of this tension in the TorC model. The right panel shows the posterior distributions of the two TorC parameters $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$ in the $H_{0}$-plane. The plot shows a positive correlation between $H_{0}$ and $\Omega_{\mathrm{L}}$, and a negative correlation between $H_{0}$ and $\varpi_{\mathrm{r}}$.
@@ IV.1 $H_{0}$ Tension
The TorC model is capable of alleviating the $H_{0}$ tension by allowing for a higher inferred Hubble parameter compared to the $\Lambda\mathrm{CDM}$ model. This is shown in the $H_{0}$-$\Omega_{\mathrm{m}}$ plane on the left panel of Fig. 5. The $\Lambda\mathrm{CDM}$ posteriors with Planck, SH0ES, and tip of the red giant branch (TRGB) [52] likelihoods reveal a clear tension between the Planck and SH0ES datasets. Whereas, the TorC posteriors for both Planck likelihood alone and joint SH0ES and Planck likelihoods show a fit that is consistent with both datasets.
The right panel of Fig. 5 shows the posterior distributions of the two TorC parameters, $\Omega_{\mathrm{L}}$ and $\varpi_{\mathrm{r}}$, along with $H_{0}$. The $H_{0}$-$\varpi_{\mathrm{r}}$ plane shows a negative correlation: as $\varpi_{\mathrm{r}}$ decreases, the value of $H_{0}$ increases. This behavior suggests that while the TorC model is consistent with $\Lambda\mathrm{CDM}$, it allows for a higher value of $H_{0}$ through the presence of torsion. Furthermore, the $H_{0}$-$\Omega_{\mathrm{L}}$ plane shows a positive correlation, indicating that a higher $H_{0}$ value in TorC is achieved not only through a reduction in $\varpi_{\mathrm{r}}$, but also by an increase in the amount of bare dark energy $\Omega_{\mathrm{L}}$.
@@ IV.3 Model comparison
Writing the logarithmic Bayes factor in the form of the second line of Eq. 23, namely $\log\Delta\mathcal{Z}=\Delta\langle\log\mathcal{L}(\theta)\rangle_{\mathcal{P}}-\Delta\mathcal{D}_{\mathrm{KL}}$, the result with Planck 2018 alone is
@@ V  Conclusions
Future work will extend the analysis of TorC parameters by incorporating other cosmological likelihoods, such as the recent DESI BAO dataset [55]. The present study assumes a vanishing torsional dark energy component, $\lambda=0$, future investigations will consider scenarios in which both bare and torsional contributions to dark energy are present. The inclusion of torsional dark energy may offer a natural explanation for the remarkably small observed value of the cosmological constant, the concept of torsion acting as a dynamical source for dark energy is also an active field of research in other extended gravity frameworks [56, 57]. This study is also conducted in the context of a flat Universe. Future work will also explore scenarios with non-zero spatial curvature. While the TorC field equations for background evolution are insensitive to the spatial curvature constant $k$, spatial curvature can still manifest through its standard geometric effects on light propagation and large-scale structure. It will be investigated whether TorC under non-zero spatial curvature can alleviate the curvature tension observed in the $\Lambda\mathrm{CDM}$ model. Finally, the development of a perturbation theory for TorC will enable a deeper exploration of the model impact on cosmic structure formation and the CMB.
@@ Acknowledgements.
The authors would like to thank Mike Hobson, Anthony Lasenby, and Dily Ong for useful discussions. SL would like to thank the IBM colour blind friendly palette for the color scheme used in the figures, and Joe Pattison and Oscar O’Hara for their advice on making the plots more visually appealing. WB is grateful for the support of Girton College, Cambridge, Marie Skłodowska-Curie Actions and the Institute of Physics of the Czech Academy of Sciences. This work used the DiRAC Data Intensive service (CSD3 www.csd3.cam.ac.uk) at the University of Cambridge, managed by the University of Cambridge University Information Services on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC component of CSD3 at Cambridge was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. The authors acknowledge the use of NumPy [58], SciPy [59], and Matplotlib [60] for numerical computations and plotting, and gemini-2.5-flash-preview-04-17 for the improvement of clarity of this paper.
@@ Appendix A Condensation and perturbation
[…]
@@
[…]
@@
[…]
@@
Figure 7: Output generated by PSALTer. The spectrograph of the theory defined by Eq. 1. At the top, the (Fourier transformed) wave operator $\mathsf{O}(k)$ from Eq. 34 is displayed in the $J^{P}$-representation defined in Tables 4 and 5. The $J^{P}$ states are organised into diagonal blocks of J=0J=0J=1J=1 and $J=2$, and within each block into diagonal parity-preserving sectors of $P=1$ (orange) and P=1P=-1 (purple), along with off-diagonal parity-violating blocks (teal). The Lagrangian in Eq. 1 does not contain any odd powers of the totally antisymmetric tensor, so there is no mixing between parities in this theory. The theory contains as many as 21 gauge generators: this far exceeds the 10 gauge generators common to all instances of Poincaré gauge theory. The gauge symmetries impose constraints on the source currents. The particle spectrum consists of two massless polarisations and one massive pseudoscalar particle (i.e. $J^{P}=0^{-}$). In terms of the couplings in Eq. 1 the pseudoscalar has mass $\sqrt{\lambda}$, so that it is not a tachyon if $\lambda\geq 0$. Its pole residue indicates that it is not a ghost if $\mu<0$. The residue of the massless pole (where $p$ denotes the particle energy in some fiducial frame) then indicates a no-ghost condition of $(\nu+2\mu)(\nu-\mu)>0$. Together, these three unitarity conditions are consistent with those in Eq. 2.
@@ Appendix B Initial conditions
$a=$ $\frac{H_{0}\sqrt{\Omega_{\mathrm{r}}}}{\alpha\varpi_{\mathrm{r}}}\tau+\frac{H_{0}^{2}\Omega_{\mathrm{m}}\bigl{(}3\varpi_{\mathrm{r}}^{2}+1\bigl{)}}{16\alpha\varpi_{\mathrm{r}}^{2}}\tau^{2}$
$+\frac{5H_{0}^{3}\Omega_{\mathrm{m}}^{2}\bigl{(}\varpi_{\mathrm{r}}^{2}-1\bigr{)}}{512\alpha\varpi_{\mathrm{r}}^{3}\sqrt{\Omega_{\mathrm{r}}}}\tau^{3}$
$+\frac{H_{0}^{4}\bigl{(}\varpi_{\mathrm{r}}^{2}-1\bigr{)}\bigl{(}27\varpi_{\mathrm{r}}^{2}-121\bigr{)}\Omega_{\mathrm{m}}^{3}}{49152\alpha\varpi_{\mathrm{r}}^{4}\Omega_{\mathrm{r}}}\tau^{4}$
$+\frac{H_{0}^{5}}{131720\alpha\varpi_{\mathrm{r}}^{5}\Omega_{\mathrm{r}}^{3/2}}\biggl{(}32768\bigl{(}3\varpi_{\mathrm{r}}^{2}+1\bigr{)}\Omega_{\mathrm{r}}^{3}\Omega_{\mathrm{L}}$
$-49\bigl{(}9\varpi_{\mathrm{r}}^{4}-29\varpi_{\mathrm{r}}^{2}+20\bigr{)}\Omega_{\mathrm{m}}^{4}\biggr{)}\tau^{5}+\mathcal{O}(\tau^{6}),$ (36a)
$\varpi=$ $\varpi_{\mathrm{r}}-\frac{3H_{0}\bigl{(}\varpi_{\mathrm{r}}^{2}-1\bigr{)}\Omega_{\mathrm{m}}}{16\sqrt{\Omega_{\mathrm{r}}}}\tau$
$+\frac{H_{0}^{2}\bigl{(}18\varpi_{\mathrm{r}}^{4}-5\varpi_{\mathrm{r}}^{2}-13\bigr{)}\Omega_{\mathrm{m}}^{2}}{512\varpi_{\mathrm{r}}\Omega_{\mathrm{r}}}\tau^{2}$
$+\frac{H_{0}^{3}\bigl{(}-324\varpi_{\mathrm{r}}^{6}+45\varpi_{\mathrm{r}}^{4}-20\varpi_{\mathrm{r}}^{2}+299\bigr{)}\Omega_{\mathrm{m}}^{3}}{49152\varpi_{\mathrm{r}}^{2}\Omega_{\mathrm{r}}^{3/2}}\tau^{3}$
$+\frac{H_{0}^{4}}{1310720\varpi_{\mathrm{r}}^{3}\Omega_{\mathrm{r}}^{2}}\biggl{[}-98304\Omega_{\mathrm{r}}^{3}\Omega_{\mathrm{L}}$
$+\bigl{(}\varpi_{\mathrm{r}}^{2}-1\bigr{)}\biggl{(}\bigl{(}2\varpi_{\mathrm{r}}^{2}\bigl{(}810(\varpi_{\mathrm{r}}^{2}+\varpi_{\mathrm{r}}^{4})+731\bigr{)}$
$+2327\bigr{)}\biggr{)}\Omega_{\mathrm{m}}^{4}\biggr{]}\tau^{4}+\mathcal{O}(\tau^{5}).$ (36b)
The parameter $\varpi_{\mathrm{r}}$ corresponds to the zeroth-order term in the power series of $\varpi$. As such, the power series of $a$ and $\varpi$ are functions of $\varpi_{\mathrm{r}}$ and $\tau$. Note that the normalisation $\alpha$ only affects the scale factor $a$. To ensure the validity of the power series, the initial value of $\tau$ used for numerical evolution is chosen to be sufficiently small such that the first term in the power series of $a$ dominates over the second term
$\tau_{\mathrm{ini}}=10^{-6}\frac{16\sqrt{\Omega_{\mathrm{r}}}\varpi_{\mathrm{r}}}{\Omega_{\mathrm{m}}\bigl{(}3\varpi_{\mathrm{r}}^{2}+1\bigr{)}H_{0}}.$ (37)
At such an early time, it is clear that the system’s evolution is primarily governed by the zeroth-order term in the power series expansion of the torsion scalar field, $\varpi_{\mathrm{r}}$.